OpenStudy (anonymous):

Im lost in trig and need a little help, (tan^4)/(sqrt(9tan^2+9)

OpenStudy (loser66):

pick denominator to solve first. what can you do ?

OpenStudy (anonymous):

could you take it out as 3tan-3?

OpenStudy (anonymous):

\[\sqrt{9\tan^2(x)+9}=\sqrt{9(\tan^2(x)+1)}=3\sqrt{\tan^2(x)+1}\] is as start

OpenStudy (anonymous):

also \(\tan^2(x)+1=\sec^2(x)\) so you can rewrite as \[\frac{\tan^4(x)}{|\sec(x)|}\]

OpenStudy (anonymous):

Then what about the top? my answer keeps coming out around \[\sin^4(x)/3\cos^2\]

OpenStudy (anonymous):

I need the answer in terms of sins and cosines,